# Joint Bayesian inference of Earth’s magnetic field and core surface flow on millennial timescales

### v1

@@ 3.3 Data model
The model likelihood is split into two parts that links the discrete time series $\mathbf{b}(t)$ to the archaeomagnetic and sedimentary palaeomagnetic data, respectively. The methodology closely follows that of [Nilsson2021], but introduces a new approach to address chronologic data uncertainties.
@@ 3.3.1 Archaeomagnetic data
$\chi_{k,i}^{(3)}=\frac{F_{k}-\textsf{H}_{k,F}(\mathbf{b}_{i})}{\sigma_{F,k}}$ (28)
@@
$P(\mathbf{o}_{k}\,|\,\mathbf{b}_{1},...,\mathbf{b}_{N},\mathbf{A},\bm{\nu})=\sum_{i=1}^{N_{t}}A_{k,i}\prod_{j=1}^{3}T_{pdf}\left(\chi_{k,i}^{(j)},\nu_{j}\right)$ (29)
where $T_{pdf}\left(\chi,\nu\right)$ is the Student’s t probability of $\chi$ with $\nu$ degrees of freedom. Note that, common to previous modelling efforts, the errors in the different field components $(D,I,F)$ are assumed to be independent, which is a reasonable assumption as long as the errors are small. The Student’s t error distribution is introduced to account for systematic uncertainties, i.e. outlier data. Following [Nilsson2021], we assign Gamma priors (table 1) for the number of degrees of freedom parameters $\bm{\nu}$, one for each magnetic field component ($D,I,F$). In practice, time steps associated low age probabilities, $A_{k,i}<0.001\Delta t/\text{yr}$ can be ignored, reducing the computational cost significantly.
@@ 3.3.2 Sedimentary data
To account for systematic directional errors we introduce the correction parameters $\delta D_{s}$ and $\delta I_{s}$ with normally distributed priors (table 1). These parameters enter equations (34) and (35) below as site-specific corrections of declination and inclination data respectively.
@@ 4 Results
To define the prior statistics for the second model, CFF.NT, we use the dynamo case 24 from [Aubert2025], previously also used and labelled as case Neutral_top1 in [Rogers2025]. This case operates at a similar level of equivalent dimensional convective power as the Coupled Earth [Aubert2013] and Midpath [Aubert2017] models, and also produces a westward-drifting columnar eccentric gyre through angular momentum conservation in the coupled outer core-inner core-mantle system. It does however use homogeneous codensity flux boundary conditions, i.e. isotropic forcing at the inner-core boundary (ICB) and the CMB, and a higher dimensional equivalent value of the electrical conductivity ($\sigma=9\times 10^{5}$ S.m-1) than those earlier models, corresponding to upward revised estimates of this parameter for Earth’s core [Pozzo2022, e.g.]. Because the dimensional rotational period of this model is about 10 years, flow variability on timescales shorter than $\sim$100 years are not considered Earth like [Gillet2024, e.g.]. For this model we therefore set $\Delta t=100$ years and $N_{v}=54$ (to retain 95% variance in the reduced rank flow vector). In addition, to facilitate sampling, the prior mean for $g_{1}^{0}$ (contained within $\mathbf{b}_{0}$) is replaced with an estimate that is consistent with the data. For the synthetic tests here we simply use the mean $g_{1}^{0}$ from Midpath.
@@ 4.1 Model-data comparison
Figure 6: Comparison of models CFF.MP based on Midpath prior (blue), CFF.NT based on Neutral_top1 (red) and the reference time series (black). From left to right: (a-f) MF $\mathbf{b}$, (e-h) total SV $\dot{\mathbf{b}}$, (i-l) induced SV $\mathbf{f}$ and (m-p) error of representativeness $\mathbf{e}$. The top panel shows the time-integrated posterior-to-prior error ratio $\Psi$ for all coefficients and below we plot time series comparisons of selected coefficients. The CFF models are represented by the posterior mean (solid lines) and 95% range (shaded areas). For a fairer comparison, the reference time series has been resampled at a 50-yr temporal resolution.
@@ 4.2 Core field and core flow
Overall, the archaeo-/palaeomagnetic data are primarily providing constraints on the toroidal flow coefficients $t_{l}^{m}$ of orders $m=0$ and $m=1$ (Fig. 7d-h). As shown in Fig. 7, the first principal component, $v_{1}$, for CFF.MP ($v_{3}$ for CFF.NT), is dominated by $t_{1}^{0}$ (Fig. 7b,e), whilst $v_{3}$ and $v_{2}$ ($v_{1}$ and $v_{2}$ in CFF.NT) mainly correspond to $t_{2}^{1,c}$ and $t_{2}^{1,s}$ (Fig. 7f,h). The latter two flow coefficients are well constrained by both models, whilst $t_{1}^{0}$ is relatively poorly resolved by CFF.NT. This is due to the statistical prior being incompatible with the reference $t_{1}^{0}$ time series in terms of variance and correlation time. The poloidal flow coefficients (Fig. 7i-k) are generally less well constrained. Notable exceptions are the $s_{1}^{1,c}$ and $s_{1}^{1,s}$ that are linked to large scale equatorial upwelling and downwelling around eastern and western limbs of the planetary-scale gyre.
@@ 4.3 Weak-field anomalies and eccentric gyres
It is not a coincident that the reference dynamo series at 1000 CE shows a weak field anomaly at Earth’s surface coinciding with eccentric gyre at roughly the same longitudes. Both the weak field anomaly and the eccentric gyre are semi-persistent features of the Midpath dynamo series that continuously drift westward and appear to be phase-locked to each other. Studying how these type of features evolve over time will be an important application of the CFF models when applied to real data. To track the weak field anomaly, we follow [Nilsson2022] and introduce a large scale field representation based exclusively on coefficients $\{g_{1}^{0},g_{2}^{1},h_{2}^{1}\}$ (Fig. 8d). Assuming a negative $g_{1}^{0}$, the amplitude (or size) of the resulting $Y_{2}^{1}$-anomaly" (East-West field asymmetry) is defined as $\sqrt{(g_{2}^{1})^{2}+(h_{2}^{1})^{2}}$ and the longitude of its centre as $\operatorname{atan2}(h_{2}^{1},g_{2}^{1})$. In a similar way we use a large scale representation of the core surface flow, based on the toroidal flow coefficients $\{t_{1}^{0},t_{2}^{1,c},t_{2}^{1,s}\}$ (Fig. 8d), to characterise the evolution of eccentric planetary gyre. Assuming a negative $t_{1}^{0}$, the amplitude (or intensity) the resulting $Y_{2}^{1}$-gyre" is defined as $\sqrt{(t_{2}^{1,c})^{2}+(t_{2}^{1,c})^{2}}$ and the longitude of the centre as $\operatorname{atan2}(t_{2}^{1,s},t_{2}^{1,c})+\pi$.
@@ 5 Discussion
Here we present the first ever integrated core field and core flow model for archaeo/palaeomagnetic MF observations and test it using synthetic data. The results demonstrate the possibility to retrieve information on large scale core surface flow structures such as the planetary scale gyre observed in modern core flow inversions [Pais2008, Gillet2019, Aubert2014, e.g.]. The reduced basis flow, as opposed to truncating the core flow expansion, allows the model to capture Earth-like core flow structures with a limited number of parameters as illustrated in Fig. 8. For the two prior dynamo series used here, the first three principal component are dominated by toroidal flow coefficients $\{t_{1}^{0},t_{2}^{1,c},t_{2}^{1,s}\}$, but it is evident from the results that higher degree and order core flow coefficients provide important contributions.
The tests highlight the influence of the dynamo prior [Suttie2025, see also]. The reliance on kinematic statistics from numerical dynamo series is both a strength and a weakness of our method. Reconstructing millennial scale variations in core flow based on sparse and noisy archaeo/palaeomagnetic observations is an inherently non-unique problem [Holme2015]. Such reconstruction will inevitably involve assumptions on (i) the role of diffusion, (ii) the influence of unmodelled sub-grid processes and (iii) the temporal and spatial characteristics of the core surface flow. Using a numerical geodynamo simulation to define the prior statistics essentially amounts to a set of physically motivated but rather strong assumptions about the dynamics of Earth’s core.
@@ Appendix A Correlated age errors
$\sigma_{DA,k}=\sigma_{a,k}\,\sigma_{p\dot{D},k}$ (38)
$\sigma_{IA,k}=\sigma_{a,k}\,\sigma_{p\dot{I},k}$ (39)
@@
In principle, $\sigma_{p\dot{D},k}^{2}$ and $\sigma_{p\dot{I},k}^{2}$ can be obtained empirically from the prior dynamo simulation. However, in cases with pDRM smoothing, this will result in an overestimation of the recorded SV variance. With increasing pDRM smoothing the correlated age uncertainties become less problematic and we therefore estimate the SV variance directly from the data using Gaussian Process regression [Rasmussen2006]. We consider the inclination data from a sediment record (after subtracting the mean and assigning ages based on the prior mean age-model, Fig. A.1a) as a realization of a zero-mean Gaussian Process we with a Matern-$\nicefrac{{3}}{{2}}$ kernel function [Gillet2013, Allington2023]
@@
Figure A.1: Tests illustrating the observational error introduced by correlated age errors using synthetic data based on the palaeomagnetic record from Gyltigesjön [Snowball2013, Nilsson2022]. The upper panel shows an example case: (a) The reference age-depth model (dashed line) drawn from the prior age-depth model (mean model shown as solid black line and 50 samples as thin gray lines). (b) The reference inclination time series $\hat{I}(t)$ (blue line) drawn from a Gaussian Process with a Matern-$\nicefrac{{3}}{{2}}$ kernel function (see text for more details). (c) Synthetic inclination data generated using the reference timeseries and reference age-depth model and additionally corrupted with Gaussian noise scaled by $\sigma_{I}$. The reference time series is projected onto the data depth-scale based on both the mean age-depth model (solid blue line) as well as the "true" reference age-depth model (dashed blue line). The lower panel: Scatter plot of log-probabilities $\log P_{T}$ and $\log P_{I}$ for 2000 cases generated as above (d) without and (e) with stratigraphic constraints. For the latter $\log P_{I}$ is evaluated using both $\sigma_{I}$ (black circles) and $\sqrt{\sigma_{I}^{2}+\sigma_{IA}^{2}}$ (red circles) as shown in (f). The case from the upper panel is shown as solid black/red circles in (e).
To illustrate the effects of ignoring stratigraphy, i.e. treating the age errors as independent, we use the palaeomagnetic record from Gyltigesjön [Snowball2013, Nilsson2022] to generate synthetic data (Fig. A.1). First, we draw 2000 random inclination time series $\hat{I}(t)$ from the Gaussian Process as defined above (Fig. A.1b). For each time series a random age-depth model (Fig. A.1a) is used to generate synthetic inclination data at the depths of measurements, corrupted by Gaussian noise according to the measurement uncertainties $\sigma_{I,k}$ (Fig. A.1c). In parallel, we construct 2000 equivalent datasets where the stratigraphy is ignored by randomly drawing the ages for each measurement from different age-depth models, resulting in truly independent age errors. For each case of $\hat{I}(t)$ we calculate the probability or likelihood (ignoring constant terms) of the $N_{GYL}$ data with known (true) ages
$\log P_{T}=\sum_{k=1}^{N_{GYL}}-\frac{\left(I_{k}-\hat{I}(t_{k})\right)^{2}}{2\sigma_{I,k}^{2}}$ (43)

### v2

@@ 3.3 Data model
[…]
@@ 3.3.1 Archaeomagnetic data
$\chi_{k,i}^{(3)}=\frac{F_{k}-\textsf{H}_{k,F}(\mathbf{b}_{i})}{\sigma_{F,k}},$ (27)
@@
$P(\mathbf{o}_{k}\,|\,\mathbf{b}_{1},...,\mathbf{b}_{N},\mathbf{A},\bm{\nu})=\sum_{i=1}^{N_{t}}A_{k,i}\prod_{j=1}^{3}T_{pdf}\left(\chi_{k,i}^{(j)},\nu_{j}\right),$ (28)
where $T_{pdf}\left(\chi,\nu\right)$ is the Student’s t probability of $\chi$ with $\nu$ degrees of freedom. Note that, common to previous modelling efforts, the errors in the different field components $(D,I,F)$ are assumed to be independent, which is a reasonable assumption as long as the errors are small. The Student’s t error distribution is introduced to account for systematic uncertainties, i.e. outlier data. Following [Nilsson2021], we assign Gamma priors (table 1) for the number of degrees of freedom parameters $\bm{\nu}$, one for each magnetic field component ($D,I,F$). In practice, time steps associated with low age probabilities, $A_{k,i}<0.001\Delta t/\text{yr}$ can be ignored, reducing the computational cost significantly. Our discrete marginalisation of the age parameters implicitly neglects variations in $\mathbf{b}(t)$ on timescales shorter than $\Delta t$. We find, however, that the associated errors (for $\Delta t<100$ yr) are negligible relative to the overall uncertainty in the data.
@@ 3.3.2 Sedimentary data
Following [Nilsson2021], to account for systematic directional errors we introduce the correction parameters $\delta D_{s}$ and $\delta I_{s}$ with normally distributed priors (table 1). These parameters enter eqs. (33) and (34) below as site-specific corrections of declination and inclination data respectively.
@@ 4 Results
To define the prior statistics for the second model, CFF.NT, we use the dynamo case 24 from [Aubert2025], previously also used and labelled as case Neutral_top1 in [Rogers2025]. This case operates at a similar level of equivalent dimensional convective power as the Coupled Earth [Aubert2013] and Midpath [Aubert2017] models, and also produces a westward-drifting columnar eccentric gyre through angular momentum conservation in the coupled outer core-inner core-mantle system. It does however use homogeneous codensity flux boundary conditions, i.e. isotropic forcing at the inner-core boundary (ICB) and the CMB, and a higher dimensional equivalent value of the electrical conductivity ($\sigma=9\times 10^{5}$ S.m-1) than those earlier models, corresponding to upward revised estimates of this parameter for Earth’s core [Pozzo2022, e.g.]. Because the dimensional rotational period of this model is about 10 years, flow variability on timescales shorter than $\sim$100 years are not considered Earth like [Gillet2024, e.g.]. For this model we therefore set $\Delta t=100$ years and $N_{v}=54$ (to retain 95% variance in the reduced rank flow vector).
[…]
@@ 4.1 Model-data comparison
Figure 6: Comparison of models CFF.MP based on Midpath prior (blue), CFF.NT based on Neutral_top1 (red) and the reference time series (black). (a-d) MF $\mathbf{b}$, (e-h) total SV $\dot{\mathbf{b}}$, (i-l) induced SV $\mathbf{f}$ and (m-p) error of representativeness $\mathbf{e}$. The top panel shows the time-integrated posterior-to-prior error ratio $\Psi$ for all coefficients and below we plot time series comparisons of selected coefficients. The CFF models are represented by the posterior mean (solid lines) and 95% range (shaded areas). For a fairer comparison, the reference time series has been resampled at a 50-yr temporal resolution.
@@ 4.2 Core field and core flow
Overall, the archaeo-/palaeomagnetic data are primarily providing constraints on the toroidal flow coefficients $t_{l}^{m}$ of orders $m=0$ and $m=1$ (Fig. 7d-h). As shown in Fig. 7, the first principal component $v_{1}$ for CFF.MP is dominated by $t_{1}^{0}$ (Fig. 7b,e), whilst $v_{3}$ and $v_{2}$ mainly correspond to $t_{2}^{1,c}$ and $t_{2}^{1,s}$ (Fig. 7f,h). The same is true for CFF.NT, but due to the different PCA basis the corresponding order of the principal components is $\{v_{3},v_{1},v_{2}\}$. The $\{t_{2}^{1,c},t_{2}^{1,s}\}$ flow coefficients are well constrained by both models, whilst $t_{1}^{0}$ is relatively poorly resolved by CFF.NT. This is due to the statistical prior being incompatible with the reference $t_{1}^{0}$ time series in terms of variance and correlation time. The poloidal flow coefficients (Fig. 7i-k) are generally less well constrained. Notable exceptions are the $s_{1}^{1,c}$ and $s_{1}^{1,s}$ that are linked to large scale equatorial upwelling and downwelling around eastern and western limbs of the planetary-scale gyre.
@@ 4.3 Weak-field anomalies and eccentric gyres
It is not a coincidence that the reference dynamo series at 1000 CE shows a weak field anomaly at Earth’s surface coinciding with eccentric gyre at roughly the same longitudes. Both the weak field anomaly and the eccentric gyre are semi-persistent features of the Midpath dynamo series that continuously drift westward and appear to be phase-locked to each other. Studying how these type of features evolve over time will be an important application of the CFF models when applied to real data. To track the weak field anomaly, we follow [Nilsson2022] and introduce a large scale field representation based exclusively on coefficients $\{g_{1}^{0},g_{2}^{1},h_{2}^{1}\}$ (Fig. 8d). Assuming a negative $g_{1}^{0}$, the amplitude (or size) of the resulting $Y_{2}^{1}$-anomaly" (East-West field asymmetry) is defined as $\sqrt{(g_{2}^{1})^{2}+(h_{2}^{1})^{2}}$ and the longitude of its centre as $\operatorname{atan2}(h_{2}^{1},g_{2}^{1})$. In a similar way we use a large scale representation of the core surface flow, based on the toroidal flow coefficients $\{t_{1}^{0},t_{2}^{1,c},t_{2}^{1,s}\}$ (Fig. 8d), to characterise the evolution of eccentric planetary gyre. Assuming a negative $t_{1}^{0}$, the amplitude (or intensity) the resulting $Y_{2}^{1}$-gyre" is defined as $\sqrt{(t_{2}^{1,c})^{2}+(t_{2}^{1,c})^{2}}$ and the longitude of the centre as $\operatorname{atan2}(t_{2}^{1,s},t_{2}^{1,c})+\pi$.
@@ 5 Discussion
Here we present the first ever integrated core field and core flow modelling framework for archaeo/palaeomagnetic MF observations and test it using synthetic data. The results demonstrate the possibility to retrieve information on large scale core surface flow structures such as the planetary scale gyre observed in modern core flow inversions [Pais2008, Gillet2019, Aubert2014, e.g.]. The reduced basis flow, as opposed to truncating the core flow expansion, allows the model to capture Earth-like core flow structures with a limited number of parameters as illustrated in Fig. 8. For the two prior dynamo series used here, the first three principal components are dominated by toroidal flow coefficients $\{t_{1}^{0},t_{2}^{1,c},t_{2}^{1,s}\}$, but it is evident from the results that higher degree and order core flow coefficients provide important contributions.
The tests highlight the influence of the dynamo prior [Suttie2025, see also]. The reliance on kinematic statistics from numerical dynamo series is both a strength and a weakness of our method. Reconstructing millennial scale variations in core flow based on sparse and noisy archaeo/palaeomagnetic observations is an inherently non-unique problem [Holme2015]. Such reconstruction will inevitably involve assumptions on (i) the role of diffusion, (ii) the influence of unmodelled sub-grid processes and (iii) the temporal and spatial characteristics of the core surface flow. Physical constraints such as quasi-geostrophy [Pais2008] can help reduce the ambiguity of core-surface flow inversions, but they do not eliminate it entirely, even for spatially well determined magnetic field [Scwaiger2023]. Given the severity of this non-uniqueness for core flow reconstructions on millennial timescales, we argue that reliance on some additional information from numerical dynamo simulations is unavoidable. In this context, using numerical geodynamo simulations to define the prior statistics represents a set of physically motivated but rather strong assumptions about the dynamics of Earth’s core.
@@ Appendix A Correlated age errors
$\sigma_{DA,k}=\sigma_{a,k}\,\sigma_{p\dot{D},k},$ (37)
$\sigma_{IA,k}=\sigma_{a,k}\,\sigma_{p\dot{I},k},$ (38)
@@
In principle, $\sigma_{p\dot{D},k}^{2}$ and $\sigma_{p\dot{I},k}^{2}$ can be obtained empirically from the prior dynamo simulation. However, in cases with pDRM smoothing, this will result in an overestimation of the recorded SV variance. With increasing pDRM smoothing the correlated age uncertainties become less problematic and we therefore estimate the SV variance directly from the data using Gaussian Process regression [Rasmussen2006]. We consider the inclination data from a sediment record (after subtracting the mean and assigning ages based on the prior mean age-model, Fig. A.1a) as a realization of a zero-mean Gaussian Process with a Matern-$\nicefrac{{3}}{{2}}$ kernel function [Gillet2013, Allington2023]
@@
Figure A.1: Tests illustrating the observational error introduced by correlated age errors using synthetic data based on the palaeomagnetic record from Gyltigesjön [Snowball2013, Nilsson2022]. The upper panel shows an example case: (a) The reference age-depth model (dashed line) drawn from the prior age-depth model (mean model shown as solid black line and 50 samples as thin gray lines). (b) The reference inclination time series $\hat{I}(t)$ (blue line) drawn from a Gaussian Process with a Matern-$\nicefrac{{3}}{{2}}$ kernel function (see text for more details). (c) Synthetic inclination data generated using the reference timeseries and reference age-depth model and additionally corrupted with Gaussian noise scaled by $\sigma_{I}$. The reference time series is projected onto the data depth-scale based on both the mean age-depth model (solid blue line) as well as the "true" reference age-depth model (dashed blue line). The lower panel: Scatter plot of log-probabilities $\log P_{T}$ and $\log P_{I}$ for 2000 cases generated as above (d) without and (e) with stratigraphic constraints. For the latter $\log P_{I}$ is evaluated using both $\sigma_{I}$ (black circles) and $\sqrt{\sigma_{I}^{2}+\sigma_{IA}^{2}}$ (red circles) as shown in (f). For illustration purposes, the $P_{T}$ and $P_{I}$ probabilities were normalised so that the respective sums of the 2000 cases equals one.
The case from the upper panel is shown as solid black/red circles in (e).
To illustrate the effects of ignoring stratigraphy, i.e. treating the age errors as independent, we use the palaeomagnetic record from Gyltigesjön [Snowball2013, Nilsson2022] to generate synthetic data (Fig. A.1). First, we draw 2000 random inclination time series $\hat{I}(t)$ from the Gaussian Process as defined above (Fig. A.1b). For each time series a random age-depth model (Fig. A.1a) is used to generate synthetic inclination data at the depths of measurements, corrupted by Gaussian noise according to the measurement uncertainties $\sigma_{I,k}$ (Fig. A.1c). In parallel, we construct 2000 equivalent datasets where the stratigraphy is ignored by randomly drawing the ages for each measurement from different age-depth models, resulting in truly independent age errors. For each case of $\hat{I}(t)$ we calculate the probability or likelihood (ignoring constant terms) of the $N_{GYL}$ data with known (true) ages:
$\log P_{T}=\sum_{k=1}^{N_{GYL}}-\frac{\left(I_{k}-\hat{I}(t_{k})\right)^{2}}{2\sigma_{I,k}^{2}},$ (42)